Anton Co. uses the perpetual inventory system and FIFO cost flow method. During the year, Anton purchased 1,040 units of inventory that cost $6 each and then purchased an additional 1,080 units of inventory that cost $8 each. If Anton sells 1,500 units of inventory, what is the amount of cost of goods sold? $12,000 $9,920 $9,000 $12,120
Answer #1

Calculation of cost of goods sold:

1040 units @6= 1040*6= 6240

460 units @ 8= 3680

Total cost of goods sold= 6240+3680= 9920

So correct answer is $9920